OpenStudy (anonymous):

Kim paid 3/7 of a bill, Sherri paid 35% of the bill, and Judy paid the rest. Which of the following statements is true? Kim paid the most. Sherri paid the most. Judy paid the most. They all paid the same amount

OpenStudy (anonymous):

Kim paid about 45% Sherri paid 35%. 45+35=78 100%-78%=22% So Kim paid the most. A is the answer

OpenStudy (harsimran_hs4):

3/7 = 0.428 % kim paid = 100 * 0.428 = 42.8 so clearly kim paid most